bill533:

Every Halloween, trick-or-treaters love going to Elise's house to see her spooky decorations and get some of her delicious caramel chews. This table shows the relationship between the number of kids in a group, x, and the total number of caramel chews Elise gives to that group, y. x (kids) y (caramel chews) 2 18 5 45 6 54 9 81 According to the values in the table, do x and y have a proportional relationship? A. yes B. no

bill533:

@563blackghost

bill533:

@Shadow

563blackghost:

To check if the table in proportional you would cross multiply the points. \(\large\bf{\frac{2}{18}=\frac{5}{45}~~~\rightarrow~~~18 \times 5=2 \times 45}\) Is the above true?

bill533:

no

bill533:

althought to let you know the numbers are separate from each other

563blackghost:

mm not quite. `18 x 5 = 90` `2 x 45 = 90` This would mean that the table `IS` proportional.

bill533:

yes

bill533:

Write an equation for the relationship between x and y. y =

bill533:

@563blackghost

silvernight269:

The answer is yes bill.